# Spin-Lattice Relaxation in Two-Dimensional Superconducting BKT Transition

### v1

@@ Abstract
Two-dimensional (2D) superconductors provide a fertile ground for studying strong phase fluctuations, where superconductivity can survive in the absence of true long-range order Lee et al. (2006); Saito et al. (2016); Qiu et al. (2021); Keimer et al. (2015). According to the Mermin–Wagner theorem, continuous symmetries cannot be spontaneously broken at finite temperature in two dimensions, ruling out conventional long-range phase coherence Hohenberg (1967); Mermin and Wagner (1966); Coleman (1973). Nevertheless, thin superconducting films and van der Waals heterostructures exhibit clear superconducting behavior, which is now understood within the framework of the Berezinskii–Kosterlitz–Thouless (BKT) transition Kosterlitz (1974); Berezinskiǐ (1972); Kosterlitz and Thouless (1973).
As schematically illustrated in Fig. 1, the superconducting transition in two dimensions is characterized by two distinct temperature scales: $T_{\rm{BCS}}$ and $T_{\rm{BKT}}$. Below $T_{\rm{BCS}}$, Cooper pairs begin to form, but strong phase fluctuations prevent the establishment of long-range order. True phase coherence is achieved only at the lower temperature $T_{\rm{BKT}}$, where a collective phase condensation occurs and the system undergoes the BKT transition. This transition is topological in nature, governed by the binding and unbinding of vortex–antivortex pairs. For $T<T_{\rm{BKT}}$, vortices remain bound, leading to quasi-long-range phase coherence and vanishing resistance. In contrast, for $T>T_{\rm{BKT}}$, vortex pairs unbind, proliferating as free vortices that destroy phase coherence and give rise to a finite resistance, even though Cooper pairs persist Ryzhov et al. (2017); Emery and Kivelson (1995); Franz and Millis (1998); Eckl et al. (2002); Eckl and Hanke (2006); Berg and Altman (2007); Han et al. (2010); Li and Han (2011); Li and Liao (2011); Li et al. (2021); Wang and Qi (2023).
Experimentally, the BKT transition is most directly identified through the universal jump of the superfluid stiffness at $T_{\mathrm{BKT}}$ Bishop and Reppy (1978); Nelson and Kosterlitz (1977); Ceperley and Pollock (1989). This criterion can be further expressed in terms of the current–voltage characteristics ($V\propto I^{\alpha}$), where the response evolves from Ohmic behavior to a non-Ohmic power law with exponent $\alpha\geq 3$ Halperin and Nelson (1979); Cotón et al. (2011); Uji et al. (2013); Zhao et al. (2013). In practice, however, the sharp IIVV jump is often obscured by disorder, inhomogeneity, and finite-size effects etc Venditti et al. (2019). This raises one question of what other experimental signatures can indicate the occurrence of the superconducting BKT transition.
@@
Since the development of Bardeen–Cooper–Schrieffer (BCS) theory, a variety of experimental probes have been employed to test the BCS pairing mechanism, including measurements of the spin-lattice relaxation rate, acoustic attenuation, and electromagnetic absorption Xiang and Wu (2022); Hebel and Slichter (1959); Morse and Bohm (1957); Tinkham (2004). In particular, the spin-lattice relaxation rate $1/T_{1}$ exhibits the well-known Hebel–Slichter (HS) coherence peak just below the critical temperature $T_{c}$, providing a key experimental confirmation of BCS pairing symmetry in conventional superconductors Rigamonti et al. (1998). Motivated by this, in this work, we investigate theoretically how the spin-lattice relaxation behaves across the BKT transition. We find that in a two-dimensional $s$-wave superconductor, a coherence peak analogous to the HS peak emerges around $T_{\rm{BKT}}$, while no clear feature appears near $T_{\rm{BCS}}$. We further extend our analysis to $d$-wave superconductors.
To capture the essential features of the BKT transition in 2D superconductivity, we consider a two-dimensional Bogoliubov–de Gennes (BdG) Hamiltonian encoding the phase fluctuations governed by an XY model Eckl et al. (2002); Han et al. (2010); Zhou et al. (2024). To simplify our discussion, we start with an s-wave superconductor. Its Hamiltonian is given by:
@@
Then, the phase $\Phi_{i}$ can be incorporated into an effective XY model and simulated using Monte Carlo methods. The free energy is $F[\Phi]=-J\sum_{ab}\text{cos}(\Phi_{a}-\Phi_{b})$, where $a$ and $b$ denote the bond between the nearest-neighbor sites for the s-wave pairing state. The parameter $J$ denotes the interaction strength in the XY model, which yields a well-established BKT transition temperature given by $T_{\rm{BKT}}=0.893J$ Komura and Okabe (2012); Nguyen and Boninsegni (2021). Since the superconducting gap $\Delta$ is typically on the order of $\Delta\sim 2T_{\rm{BCS}}$, we choose $J=0.112$ such that $T_{\rm{BKT}}=0.1$, ensuring that it lies well below the mean-field transition temperature $T_{\rm{BCS}}$. The Monte Carlo simulations are performed on a $N=72\times 72$ square lattice. To further suppress finite-size effects, we employed twisted boundary conditions (see details in the Supplementary Materials).
The BKT transition is topological rather than thermodynamic in nature; hence, its transition point cannot be characterized by singularities in specific heat. In Fig. 2(a), we show the temperature dependence of the superfluid density. Blue squares represent the superfluid density of bosons in the XY model Canova et al. (2014), while red circles show the fermionic superfluid density Scalapino et al. (1992, 1993) computed from the BdG Hamiltonian. The BKT transition temperature is determined by the intersection of the superfluid density curve and the universal line described by the scaling relation $\rho_{s}(T)/\rho_{s}(0)\sim 2T/\pi J$, denoted by the black dashed dot line in Fig. 2 (a). Due to the discrete energy spectrum inherent to a finite fermionic lattice, the numerically computed superfluid density for fermions exhibits deviations from that of the effective bosonic model. Nevertheless, the intersection point with the scaling line remains consistent. Since the phase fluctuations in our model are governed by an XY Hamiltonian, the bosonic phase field $\{\Phi_{i}\}$ directly controls vortex formation, pairing, and decoupling processes. Therefore, throughout this work, we use the bosonic superfluid density to locate the BKT transition.
For each configuration of the phase field $\Phi$}, the order parameter $\Delta_{i}$ varies from site to site, breaking the translational invariance of the system. Under such conditions, the spin-lattice relaxation rate is related to magnetic susceptibility $\chi_{zz}$ via Xiang and Wu (2022); Slichter (2013):
@@
where $j$ and $l$ denote the coordinates of site $i$ or its neighboring lattice points. Here, $F_{j,i}$ represents the structure factor of the hyperfine interaction Slichter (2013). To reduce computational cost, we include only the dominant contribution from the term with $j=l=i$. In the limit $\omega\to 0$, the magnetic susceptibility $\operatorname{Im}\chi_{zz}(j,l,\omega)$ can be expressed using the electron Green’s function as: $\lim_{\omega\to 0}\frac{\operatorname{Im}\chi_{zz}(j,l,\omega)}{\omega}=-\frac{\gamma_{e}^{2}\hbar^{2}}{2\pi}\int_{-\infty}^{\infty}\mathrm{d}\varepsilon A(j,l,\varepsilon)\frac{\partial f(\varepsilon)}{\partial\varepsilon}.$ Here $A(j,j^{\prime},\varepsilon)=\left[\operatorname{Im}G_{11}(j,j^{\prime};\varepsilon)\right]^{2}+\left[\operatorname{Im}G_{12}(j,j^{\prime};\varepsilon)\right]^{2}$, where the Green’s functions $G_{11}$ and $G_{12}$ represent the normal and anomalous Green’s function, respectively. Thus $1/T_{1}T$ can be further rewritten as:
@@
Fig. 2 (b) shows the calculated spin-lattice relaxation rate $1/T_{1}T$. At high temperatures, above roughly $2T_{\rm{BKT}}$, $1/T_{1}T$ remains nearly constant. As the temperature approaches $T_{\rm{BKT}}$, a gradual increase develops, culminating in a pronounced peak near the BKT transition. This peak closely resembles the HS coherence peak observed in conventional superconductors MacLaughlin (1976); Tinkham (2004); Rigamonti et al. (1998); Schrieffer (2018); Hebel and Slichter (1957, 1959); Hebel (1959). Since this system remains a disordered system, the peak is broader than BCS theory. At lower temperatures, $1/T_{1}T$ decreases gradually and vanishes as the system enters the phase-coherent superconducting state.
Physically, the origin of the HS peak can be traced to the singular behavior of the quasiparticle DOS in BCS theory. When the superconducting gap opens below $T_{c}$, the Bogoliubov quasiparticle spectrum produces a square-root divergence at the gap edge Schrieffer (2018). This coherence-enhanced DOS strongly amplifies the available scattering channels for nuclear spins, thereby increasing the spin-lattice relaxation rate just below the transition Slichter (2013). The enhancement is further reinforced by the BCS coherence factors, which govern the quasiparticle matrix elements entering the relaxation process Schrieffer (2018). Together, these effects generate the characteristic HS peak that serves as a hallmark of conventional $s$-wave superconductivity Schrieffer (2018). At lower temperatures, however, the number of thermally excited quasiparticles decreases rapidly as the superconducting gap fully develops. Since the relaxation rate is proportional to the quasiparticle population, $1/T_{1}T$ drops and vanishes in the zero-temperature limit, reflecting the absence of low-energy excitations in a fully gapped $s$-wave state.
@@
Hence, we can conclude that there is an HS-like peak in $1/T_{1}T$ during the 2D superconducting BKT transition. This also originates from the DOS singularity occurring around $T_{\rm{BKT}}$. Furthermore, we need to check the $1/T_{1}T$ feature around $T_{\rm{BCS}}$ to confirm our results. Then, we can include the gap amplitude fluctuation through a BCS self-consistent gap equation gap . To focus on the $T_{\rm{BCS}}$ region and ensure that the BKT transition temperature lies well below $T_{\rm{BCS}}$, we adopt a reduced XY-model interaction corresponding to $T_{\rm{BKT}}=0.04$. The gap function $|\Delta|(T)$ with $T_{\rm{BCS}}=0.17$ is plotted in Fig. 3(a). Below $T_{\rm{BCS}}$, the $|\Delta|(T)$ becomes finite and saturates at low temperature following a standard BCS feature. Correspondingly, $1/T_{1}T$ is also calculated within the same framework, shown in Fig. 3(b). We can find that the $1/T_{1}T$ shows no peak or any distinct feature crossing $T_{\rm{BCS}}$. This confirms that the coherence peak observed in our model is exclusively a signature of the BKT transition, driven by the establishment of long-range phase coherence, and is not associated with the formation of Cooper pairs at $T_{\rm{BCS}}$.
$d$-wave We can also extend our study to $d$-wave superconductors and explore the $1/T_{1}T$ response crossing $T_{\rm{BKT}}$. It is well-established that no HS peaks are observed in cuprates Rigamonti et al. (1998). The pairing term for $d$-wave SC of the Hamiltonian Eq. (1) is expressed as:
@@
To further understand the discrepancy of the spin-lattice relaxation rate between the cases of $s$-wave and $d$-wave pairing, we plot the DOS at several temperatures in Fig. 4 (c). In contrast to the s-wave case, the DOS exhibits a stable V-shaped gap. For $T<T_{\rm{BKT}}$, coherence peaks emerge and their amplitude increases with decreasing temperature. The absence of HS peak, which instead emerges in its derivation with temperature, can be understood by analyzing three distinct temperature regimes. Below $T_{\rm{BKT}}$, there exists only a smaller spectral contribution near the coherence peaks, and the zero-frequency DOS remains almost unchanged; thus, the slight increase in $1/T_{1}T$ with increasing $T$ is mainly due to thermal effects. Above $T_{\rm{BKT}}$, the coherence peaks vanish completely. The increase in the in-gap spectral weight leads to a gradual increase in $1/T_{1}T$. Once the pseudogap is fully suppressed, $1/T_{1}T$ approaches a constant value, consistent with the behavior of a conventional metal at high temperatures. In the high-$T$ regime ($T>T_{\rm{BKT}}$), the temperature dependence of $1/T_{1}T$ fully correlates with that of $A_{\omega}(\omega=0)$. Crucially, this correspondence extends to their respective temperature derivatives. As shown in Fig. 4 (d), there also exists a peak around $T\sim 0.118$. The zero-frequency DOS $A_{\omega}(\omega=0)$ and its derivative $\frac{\partial A_{\omega}(\omega=0)}{\partial T}$ reveal a gradual emergence and an increase of in-gap spectral weight within the pseudogap regime for $T_{\rm BKT}<T<T_{\rm BCS}$. This theoretical result is strikingly similar to experimental observations in the cuprate superconductor Bi2212Chen et al. (2022). Although its transition is not of the BKT type, Bi2212 is also characterized by pre-formed pairs above $T_{c}$ whose coherence is destroyed by strong phase fluctuations. Critically, a peak in $\frac{\partial A_{\omega}(\omega=0)}{\partial T}$ is also observed experimentally at $T_{c}$ in this material, reinforcing its interpretation as a clear signature of the onset of macroscopic phase coherence.
@@
Acknowledgments We would like to thank Yin Zhong for useful discussions. We acknowledge the support by the Ministry of Science and Technology (Grant No. 2022YFA1403900), the National Natural Science Foundation of China (Grant NSFC-12494594, No. NSFC-12174428, No. 12474146, No. 12347107), the New Cornerstone Investigator Program, the Chinese Academy of Sciences Project for Young Scientists in Basic Research (2022YSBR-048), and the Beijing Natural Science Foundation (No. JR25007).
@@
Our numerical approach is a hybrid method that combines a classical treatment of order parameter fluctuations with an exact quantum description of the fermions. The thermal fluctuations of the superconducting phase are modeled using a long-wavelength XY model and sampled via standard Monte Carlo (MC) simulations. For each classical phase configuration generated, we determine the single-particle fermionic spectrum by numerically diagonalizing the corresponding quadratic Hamiltonian. While the repeated diagonalization of the fermionic Hamiltonian constitutes the primary computational bottleneck, this hybrid scheme is highly efficient, enabling the study of large systems up to $N=72\times 72$ lattice sites. To minimize finite-size effects, we implement twisted boundary conditions by threading a magnetic flux through the lattice for both spin sectors. This is accomplished with a Peierls substitution in the Landau gauge, $\mathbf{A}=(-By,0,0)$ Assaad (2002); Schattner et al. (2016), which modifies the nearest-neighbor hopping terms such that the hopping in the $\hat{x}-$direction acquires a position-dependent phase factor:
@@
To investigate its origin, we plot the waterfall diagram of the local density of states along a horizontal cut across the lattice at various temperatures. As illustrated in Fig. S1, the absence of translational symmetry leads to spatial variations in the gap magnitude. As temperature increases, the gap size decreases. Near $T_{\rm{BKT}}$, the coherence peaks at the gap edges shift toward zero energy, significantly influencing $1/T_{1}T$. Below $T_{\rm{BKT}}$, vortex pairs remain confined due to the BKT transition, permitting a finite gap on every site. However, as temperature rises further to $T\gtrsim T_{\rm{BKT}}$, proliferation of unbound vortices is expected. The presence of free vortices forces the gap magnitude to vanish at the vortex cores, i.e., $\Delta_{i}=0$. This gives rise to finite subgap density of states with small peaks at these sites. Combined with the persistence of coherence peaks at the gap edges, these states further enhance $1/T_{1}T$, resulting in the observed NMR peak.
Figure S2 presents the waterfall plot of the local density of states for the d-wave superconductor. Combined with Fig. 3(b), these results reveal two distinct regimes in the DOS behavior. First, in the superconducting regime ($T<T_{\rm{BKT}}$), the coherence peak amplitude at the gap edges gradually decreases with increasing temperature, where even the zero-temperature coherence peak amplitude is significantly reduced compared to that in the s-wave case. Second, in the pseudogap regime ($T_{\rm{BKT}}<T<2T_{\rm{BKT}}$), the in-gap spectral weight increases steadily until the pseudogap is completely filled.
Figure S1: Waterfall plot of the density of states for s-wave pairing at various temperatures. Each curve represents the local density of states along a horizontal cut across the lattice. As temperature increases, the superconducting gap gradually decreases. Bound states emerge near $T_{c}$, as highlighted by the red dashed box. Above $T_{c}$, the gap is smeared out by thermal effects: the U-shaped gap closes and the coherence peak characteristic of s-wave pairing disappears.
Figure S2: Waterfall plot of the density of states for d-wave pairing at various temperatures. Each curve corresponds to the local density of states along a horizontal cut across the sample. In contrast to s-wave pairing, the density of states near zero energy remains largely unchanged with increasing temperature. However, the coherence peaks gradually diminish below $T_{c}$ and are strongly suppressed near and above $T_{c}$. The combined absence of clear bound states and the suppression of coherence peaks result in no distinctive features in the density of states that would contribute to the spin–lattice relaxation rate, leading to its monotonic increase with temperature. Above $T_{c}$ the density of states is smeared by thermal fluctuations.

### v2

@@ Abstract
Two-dimensional (2D) superconductors provide a fertile ground for studying strong phase fluctuations, where superconductivity can survive in the absence of true long-range order RevModPhys.78.17; Saito2016; qiu2021recent; keimer2015quantum. According to the Mermin–Wagner theorem, continuous symmetries cannot be spontaneously broken at finite temperature in two dimensions, ruling out conventional long-range phase coherence PhysRev.158.383; PhysRevLett.17.1133; Coleman1973. Nevertheless, thin superconducting films and van der Waals heterostructures exhibit clear superconducting behavior, which is now understood within the framework of the Berezinskii–Kosterlitz–Thouless (BKT) transition J_M_Kosterlitz_1974; 1972JETP_34..610B; J_M_Kosterlitz_1973.
As schematically illustrated in Fig. 1, the superconducting transition in two dimensions is characterized by two distinct temperature scales: $T_{\rm{BCS}}$ and $T_{\rm{BKT}}$. Below $T_{\rm{BCS}}$, Cooper pairs begin to form, but strong phase fluctuations prevent the establishment of long-range order. True phase coherence is achieved only at the lower temperature $T_{\rm{BKT}}$, where a collective phase condensation occurs and the system undergoes the BKT transition. This transition is topological in nature, governed by the binding and unbinding of vortex–antivortex pairs. For $T<T_{\rm{BKT}}$, vortices remain bound, leading to quasi-long-range phase coherence and vanishing resistance. In contrast, for $T>T_{\rm{BKT}}$, vortex pairs unbind, proliferating as free vortices that destroy phase coherence and give rise to a finite resistance, even though Cooper pairs persist Ryzhov:2017; Emery1995; PhysRevB.58.14572; PhysRevB.66.140510; PhysRevB.74.134510; PhysRevLett.99.247001; PhysRevB.82.052503; Li_2011; Li_2011b; li2021superconductor; PhysRevB.107.224502.
Experimentally, the BKT transition is most directly identified through the universal jump of the superfluid stiffness at $T_{\mathrm{BKT}}$ PhysRevLett.40.1727; PhysRevLett.39.1201; PhysRevB.39.2084. This criterion can be further expressed in terms of the current–voltage characteristics ($V\propto I^{\alpha}$), where the response evolves from Ohmic behavior to a non-Ohmic power law with exponent $\alpha\geq 3$ Halperin; Cotón_2011; PhysRevLett.110.196602; ZHAO201359. In practice, however, the sharp IIVV jump is often obscured by disorder, inhomogeneity, and finite-size effects etc PhysRevB.100.064506. This raises one question of what other experimental signatures can indicate the occurrence of the superconducting BKT transition.
@@
Since the development of Bardeen–Cooper–Schrieffer (BCS) theory, a variety of experimental probes have been employed to test the BCS pairing mechanism, including measurements of the spin-lattice relaxation rate, acoustic attenuation, and electromagnetic absorption Xiang2022; PhysRev.113.1504; PhysRev.108.1094; tinkham2004introduction. In particular, the spin-lattice relaxation rate $1/T_{1}$ exhibits the well-known Hebel–Slichter (HS) coherence peak just below the critical temperature $T_{c}$, providing a key experimental confirmation of BCS pairing symmetry in conventional superconductors A_Rigamonti_1998. Motivated by this, in this work, we investigate theoretically how the spin-lattice relaxation behaves across the BKT transition. We find that in a two-dimensional $s$-wave superconductor, a coherence peak analogous to the HS peak emerges around $T_{\rm{BKT}}$, while no clear feature appears near $T_{\rm{BCS}}$. We further extend our analysis to $d$-wave superconductors.
To capture the essential features of the BKT transition in 2D superconductivity, we consider a two-dimensional Bogoliubov–de Gennes (BdG) Hamiltonian encoding the phase fluctuations governed by an XY model PhysRevB.66.140510; PhysRevB.82.052503; zhou2024universalscalingbehaviorresistivity. To simplify our discussion, we start with an s-wave superconductor. Its Hamiltonian is given by:
@@
Then, the phase $\Phi_{i}$ can be incorporated into an effective XY model and simulated using Monte Carlo methods. The free energy is $F[\Phi]=-J\sum_{ab}\text{cos}(\Phi_{a}-\Phi_{b})$, where $a$ and $b$ denote the bond between the nearest-neighbor sites for the s-wave pairing state. The parameter $J$ denotes the interaction strength in the XY model, which yields a well-established BKT transition temperature given by $T_{\rm{BKT}}=0.893J$ doi:10.1143/JPSJ.81.113001; app11114931. Since the superconducting gap $\Delta$ is typically on the order of $\Delta\sim 2T_{\rm{BCS}}$, we choose $J=0.112$ such that $T_{\rm{BKT}}=0.1$, ensuring that it lies well below the mean-field transition temperature $T_{\rm{BCS}}$. The Monte Carlo simulations are performed on a $N=72\times 72$ square lattice. To further suppress finite-size effects, we employed twisted boundary conditions (see details in the Supplementary Materials).
The BKT transition is topological rather than thermodynamic in nature; hence, its transition point cannot be characterized by singularities in specific heat. In Fig. 2(a), we show the temperature dependence of the superfluid density. Blue squares represent the superfluid density of bosons in the XY model Canova_2014, while red circles show the fermionic superfluid density PhysRevLett.68.2830; PhysRevB.47.7995 computed from the BdG Hamiltonian. The BKT transition temperature is determined by the intersection of the superfluid density curve and the universal line described by the scaling relation $\rho_{s}(T)/\rho_{s}(0)\sim 2T/\pi J$, denoted by the black dashed dot line in Fig. 2 (a). Due to the discrete energy spectrum inherent to a finite fermionic lattice, the numerically computed superfluid density for fermions exhibits deviations from that of the effective bosonic model. Nevertheless, the intersection point with the scaling line remains consistent. Since the phase fluctuations in our model are governed by an XY Hamiltonian, the bosonic phase field $\{\Phi_{i}\}$ directly controls vortex formation, pairing, and decoupling processes. Therefore, throughout this work, we use the bosonic superfluid density to locate the BKT transition.
For each configuration of the phase field $\Phi$}, the order parameter $\Delta_{i}$ varies from site to site, breaking the translational invariance of the system. Under such conditions, the spin-lattice relaxation rate is related to magnetic susceptibility $\chi_{zz}$ via Xiang2022; slichter2013principles:
@@
where $j$ and $l$ denote the coordinates of site $i$ or its neighboring lattice points. Here, $F_{j,i}$ represents the structure factor of the hyperfine interaction slichter2013principles. To reduce computational cost, we include only the dominant contribution from the term with $j=l=i$. In the limit $\omega\to 0$, the magnetic susceptibility $\operatorname{Im}\chi_{zz}(j,l,\omega)$ can be expressed using the electron Green’s function as: $\lim_{\omega\to 0}\frac{\operatorname{Im}\chi_{zz}(j,l,\omega)}{\omega}=-\frac{\gamma_{e}^{2}\hbar^{2}}{2\pi}\int_{-\infty}^{\infty}\mathrm{d}\varepsilon A(j,l,\varepsilon)\frac{\partial f(\varepsilon)}{\partial\varepsilon}.$ Here $A(j,j^{\prime},\varepsilon)=\left[\operatorname{Im}G_{11}(j,j^{\prime};\varepsilon)\right]^{2}+\left[\operatorname{Im}G_{12}(j,j^{\prime};\varepsilon)\right]^{2}$, where the Green’s functions $G_{11}$ and $G_{12}$ represent the normal and anomalous Green’s function, respectively. Thus $1/T_{1}T$ can be further rewritten as:
@@
Fig. 2 (b) shows the calculated spin-lattice relaxation rate $1/T_{1}T$. At high temperatures, above roughly $2T_{\rm{BKT}}$, $1/T_{1}T$ remains nearly constant. As the temperature approaches $T_{\rm{BKT}}$, a gradual increase develops, culminating in a pronounced peak near the BKT transition. This peak closely resembles the HS coherence peak observed in conventional superconductors MACLAUGHLIN19761; tinkham2004introduction; A_Rigamonti_1998; schrieffer2018theory; PhysRev.107.901; PhysRev.113.1504; PhysRev.116.79. Since this system remains a disordered system, the peak is broader than BCS theory. At lower temperatures, $1/T_{1}T$ decreases gradually and vanishes as the system enters the phase-coherent superconducting state.
Physically, the origin of the HS peak can be traced to the singular behavior of the quasiparticle DOS in BCS theory. When the superconducting gap opens below $T_{c}$, the Bogoliubov quasiparticle spectrum produces a square-root divergence at the gap edge schrieffer2018theory. This coherence-enhanced DOS strongly amplifies the available scattering channels for nuclear spins, thereby increasing the spin-lattice relaxation rate just below the transition slichter2013principles. The enhancement is further reinforced by the BCS coherence factors, which govern the quasiparticle matrix elements entering the relaxation process schrieffer2018theory. Together, these effects generate the characteristic HS peak that serves as a hallmark of conventional $s$-wave superconductivity schrieffer2018theory. At lower temperatures, however, the number of thermally excited quasiparticles decreases rapidly as the superconducting gap fully develops. Since the relaxation rate is proportional to the quasiparticle population, $1/T_{1}T$ drops and vanishes in the zero-temperature limit, reflecting the absence of low-energy excitations in a fully gapped $s$-wave state.
@@
Hence, we can conclude that there is an HS-like peak in $1/T_{1}T$ during the 2D superconducting BKT transition. This also originates from the DOS singularity occurring around $T_{\rm{BKT}}$. Furthermore, we need to check the $1/T_{1}T$ feature around $T_{\rm{BCS}}$ to confirm our results. Then, we can include the gap amplitude fluctuation through a BCS self-consistent gap equation gap_function. To focus on the $T_{\rm{BCS}}$ region and ensure that the BKT transition temperature lies well below $T_{\rm{BCS}}$, we adopt a reduced XY-model interaction corresponding to $T_{\rm{BKT}}=0.04$. The gap function $|\Delta|(T)$ with $T_{\rm{BCS}}=0.17$ is plotted in Fig. 3(a). Below $T_{\rm{BCS}}$, the $|\Delta|(T)$ becomes finite and saturates at low temperature following a standard BCS feature. Correspondingly, $1/T_{1}T$ is also calculated within the same framework, shown in Fig. 3(b). We can find that the $1/T_{1}T$ shows no peak or any distinct feature crossing $T_{\rm{BCS}}$. This confirms that the coherence peak observed in our model is exclusively a signature of the BKT transition, driven by the establishment of long-range phase coherence, and is not associated with the formation of Cooper pairs at $T_{\rm{BCS}}$. In contrast to the featureless behavior of $1/T_{1}T$, the Knight shift exhibits a distinct kink at $T_{\rm{BCS}}$ for both $s$- and $d$-wave symmetries, followed by a linear decrease confined to the region immediately below $T_{\rm{BCS}}$ supp_knight.
$d$-wave We can also extend our study to $d$-wave superconductors and explore the $1/T_{1}T$ response crossing $T_{\rm{BKT}}$. It is well-established that no HS peaks are observed in cuprates A_Rigamonti_1998. The pairing term for $d$-wave SC of the Hamiltonian Eq. (1) is expressed as:
@@
To further understand the discrepancy of the spin-lattice relaxation rate between the cases of $s$-wave and $d$-wave pairing, we plot the DOS at several temperatures in Fig. 4 (c). In contrast to the s-wave case, the DOS exhibits a stable V-shaped gap. For $T<T_{\rm{BKT}}$, coherence peaks emerge and their amplitude increases with decreasing temperature. The absence of HS peak, which instead emerges in its derivation with temperature, can be understood by analyzing three distinct temperature regimes. Below $T_{\rm{BKT}}$, there exists only a smaller spectral contribution near the coherence peaks, and the zero-frequency DOS remains almost unchanged; thus, the slight increase in $1/T_{1}T$ with increasing $T$ is mainly due to thermal effects. Above $T_{\rm{BKT}}$, the coherence peaks vanish completely. The increase in the in-gap spectral weight leads to a gradual increase in $1/T_{1}T$. Once the pseudogap is fully suppressed, $1/T_{1}T$ approaches a constant value, consistent with the behavior of a conventional metal at high temperatures. In the high-$T$ regime ($T>T_{\rm{BKT}}$), the temperature dependence of $1/T_{1}T$ fully correlates with that of $A_{\omega}(\omega=0)$. Crucially, this correspondence extends to their respective temperature derivatives. As shown in Fig. 4 (d), there also exists a peak around $T\sim 0.118$. The zero-frequency DOS $A_{\omega}(\omega=0)$ and its derivative $\frac{\partial A_{\omega}(\omega=0)}{\partial T}$ reveal a gradual emergence and an increase of in-gap spectral weight within the pseudogap regime for $T_{\rm BKT}<T<T_{\rm BCS}$. This theoretical result is strikingly similar to experimental observations in the cuprate superconductor Bi2212Chen2022. Although its transition is not of the BKT type, Bi2212 is also characterized by pre-formed pairs above $T_{c}$ whose coherence is destroyed by strong phase fluctuations. Critically, a peak in $\frac{\partial A_{\omega}(\omega=0)}{\partial T}$ is also observed experimentally at $T_{c}$ in this material, reinforcing its interpretation as a clear signature of the onset of macroscopic phase coherence.
@@
Acknowledgments We would like to thank Yin Zhong for useful discussions. We acknowledge the support by the Ministry of Science and Technology (Grant No. 2022YFA1403900), the National Natural Science Foundation of China (Grant NSFC-12494594, No. NSFC-12174428, No. 12474146, No. 12347107), the New Cornerstone Investigator Program, the Chinese Academy of Sciences Project for Young Scientists in Basic Research (2022YSBR-048), and the Beijing Natural Science Foundation (No. JR25007). W. Y. acknowledge support from the China Postdoctoral Science Foundation (Grant No. 2025M773414).
@@
Our numerical approach is a hybrid method that combines a classical treatment of order parameter fluctuations with an exact quantum description of the fermions. The thermal fluctuations of the superconducting phase are modeled using a long-wavelength XY model and sampled via standard Monte Carlo (MC) simulations. For each classical phase configuration generated, we determine the single-particle fermionic spectrum by numerically diagonalizing the corresponding quadratic Hamiltonian. While the repeated diagonalization of the fermionic Hamiltonian constitutes the primary computational bottleneck, this hybrid scheme is highly efficient, enabling the study of large systems up to $N=72\times 72$ lattice sites. To minimize finite-size effects, we implement twisted boundary conditions by threading a magnetic flux through the lattice for both spin sectors. This is accomplished with a Peierls substitution in the Landau gauge, $\mathbf{A}=(-By,0,0)$ PhysRevB.65.115104; PhysRevLett.117.097002, which modifies the nearest-neighbor hopping terms such that the hopping in the $\hat{x}-$direction acquires a position-dependent phase factor:
@@
To investigate its origin, we plot the waterfall diagram of the local density of states (DOS) along a horizontal cut across the lattice at various temperatures. As illustrated in Fig. S1, the absence of translational symmetry leads to spatial variations in the gap magnitude. As temperature increases, the gap size decreases. Near $T_{\rm{BKT}}$, the coherence peaks at the gap edges shift toward zero energy, significantly influencing $1/T_{1}T$. Below $T_{\rm{BKT}}$, vortex pairs remain confined due to the BKT transition, permitting a finite gap on every site. However, as temperature rises further to $T\gtrsim T_{\rm{BKT}}$, proliferation of unbound vortices is expected. The presence of free vortices forces the gap magnitude to vanish at the vortex cores, i.e., $\Delta_{i}=0$. This gives rise to finite subgap DOS with small peaks at these sites. Combined with the persistence of coherence peaks at the gap edges, these states further enhance $1/T_{1}T$, resulting in the observed NMR peak.
Figure S2 presents the waterfall plot of the local DOS for the d-wave superconductor. Combined with Fig. 3(b), these results reveal two distinct regimes in the DOS behavior. First, in the superconducting regime ($T<T_{\rm{BKT}}$), the coherence peak amplitude at the gap edges gradually decreases with increasing temperature, where even the zero-temperature coherence peak amplitude is significantly reduced compared to that in the s-wave case. Second, in the pseudogap regime ($T_{\rm{BKT}}<T<2T_{\rm{BKT}}$), the in-gap spectral weight increases steadily until the pseudogap is completely filled.
Figure S1: Waterfall plot of the DOS for s-wave pairing at various temperatures. Each curve represents the local DOS along a horizontal cut across the lattice. As temperature increases, the superconducting gap gradually decreases. Bound states emerge near $T_{c}$, as highlighted by the red dashed box. Above $T_{c}$, the gap is smeared out by thermal effects: the U-shaped gap closes and the coherence peak characteristic of s-wave pairing disappears.
Figure S2: Waterfall plot of the DOS for d-wave pairing at various temperatures. Each curve corresponds to the local DOS along a horizontal cut across the sample. In contrast to s-wave pairing, the DOS near zero energy remains largely unchanged with increasing temperature. However, the coherence peaks gradually diminish below $T_{c}$ and are strongly suppressed near and above $T_{c}$. The combined absence of clear bound states and the suppression of coherence peaks result in no distinctive features in the DOS that would contribute to the spin–lattice relaxation rate, leading to its monotonic increase with temperature. Above $T_{c}$ the DOS is smeared by thermal fluctuations.